OpenStudy (anonymous):

Sherri has 23 pieces of jewelry to sell. She sells the bracelets for $2 and the necklaces for $3, and earns a total of $55. If the bracelets are represented by x and the necklaces are represented by y, which of the following systems of equations can be used to calculate the number of bracelets and necklaces sold?

OpenStudy (anonymous):

x+y=23, 3x+2y = 55 x + y = 55, 3x + 2y = 55 x+y=23, 2x+3y= 55 x+y=55, 2x+3y =23

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

c

OpenStudy (anonymous):

@waleedo212 can you explain why?

OpenStudy (anonymous):

so they said the sum is x+y so if the sum is 55, x+y=55

OpenStudy (anonymous):

sorry im eant x+y =23 pieces

OpenStudy (anonymous):

secondly, 2x+3y, represents the cost of the pieces, and since they mentioned the amount of the sale is 55, it is 2x+3y=55

OpenStudy (anonymous):

Okay that makes sense. Thanks!

OpenStudy (anonymous):

np :P